OpenStudy (yb1996):

find the limit as x approaches infinity of (x^6+18x^4)^(1/3)-x^2

OpenStudy (yb1996):

The correct answer is 6. I just don't know how to get that.

OpenStudy (anonymous):

Factor \(x^6\) out of the radicand,\[\lim_{x \rightarrow \infty} \sqrt[3]{x^6\left( 1+\frac{ 18 }{ x^2 } \right)} - x^2\]

OpenStudy (yb1996):

Then I get \[\lim_{x \rightarrow \infty} x^2\sqrt[3]{1+18/x^2}-x^2\]

OpenStudy (anonymous):

OK What happens when you factor out the common factor of x^2? (Hope I'm not leading you down the garden path)

OpenStudy (anonymous):

You're right, the correct answer is 6. Perhaps my thoughts are leading nowhere.

OpenStudy (yb1996):

That would give me \[x^2(\sqrt[3]{\frac{ 18 }{ x^2 }+1}-1)\]

OpenStudy (yb1996):

which means as x goes to infinity, the limit will also be going to infinity, which is not right

OpenStudy (anonymous):

Right-o. I was looking at the quantity in brackets going to zero.

OpenStudy (yb1996):

Oh, never mind. So it's infinity times zero.

OpenStudy (yb1996):

But how do you continue the problem?

OpenStudy (anonymous):

Thinking...

OpenStudy (anonymous):

A series expansion of that expression looks like\[6-\frac{ 36 }{ x^2 }+\frac{ 360 }{ x^4 }-\frac{ 4320 }{ x^6 }+...\]which shows how the limit approaches 6. But there must be an easier way.

OpenStudy (yb1996):

Wait, how did you do that?

OpenStudy (anonymous):

It's a Laurent series expansion. I looked it up. But it's a long time since I've done them. You'd be better off researching doing Laurent expansions. I'd just mess it up trying to remember how to do it.

OpenStudy (yb1996):

I did it! Thanks for starting me off, it actually helped!!!

OpenStudy (anonymous):

Good for you! Job well done